OpenStudy (anonymous):

Find the area of the triangle shown on the right with base 1 2/3 and height 3/7 yard (hint the area of a triangle equals half the product of it's base and height)

OpenStudy (anonymous):

This one is a give me. You simply plug in the values for Area=1/2bh where b=5/3 and h=3/7.... so you get (1/2)(5/3)(3/7)=5/14units^2